Determine the new concentration of Zn^2+ ion in gr


Help.
Answer #1

i) molarity of mixer = M1V1+M2V2/V1+V2

   = (50*0.025+100*0.15)/150

   = 0.1083 M

molarity of Zn^2+ = 0.1083 M

ii)

50 ml = 0.5 dl

150 micrograms = 0.15 mg

concentration of milk = 0.15/0.5 = 0.3 mg/dl

i)

from standard solution

90 grams HNO3 present in 100 grams solution

volume of solution = m/density = 100/1.51 = 66.22 ml

Molarity of standard solution = (90/63)*(1000/66.22) = 21.57 M

dilution formula

M1V1 = M2V2

21.57*V1 = 0.25*500

V1 = 5.8 ml

take 5.8 ml standard solution and add water up to 500 ml

ii)

150 ppm means

150 mg present in 1 liter solution

Molarity of OH- = (0.15/17)*(1/1) = 0.0088 M

1 mol Ba(OH)2 = 2 mol OH-

No of mol of Ba(OH)2 required = 0.0088/2 = 0.0044 mol

mass of Ba(OH)2 MUST DISSOLVE = 0.0044*171.34 = 0.754 grams